# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@
The Very High Energy Afterglow of Structured Jets:
@@
Clément Pellouin, The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections    Frédéric Daigne, The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future DetectionsThe Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections
@@
We present a complete numerical model of the afterglow of a laterally-structured relativistic ejecta from radio to very high energy (VHE). This includes a self-consistent calculation of the synchrotron radiation, with its maximum frequency, and of Synchrotron Self-Compton (SSC) scatterings taking into account the Klein-Nishina regime. The attenuation due to pair production is also included. This model is computationally-efficient, allowing for multi-wavelength data fitting. As a validation test, the radiative model is used to fit the broad-band spectrum of GRB 190114C at 90 s up to the TeV range. The full model is then used to fit the afterglow of GW 170817 and predict its VHE emission. We find that the SSC flux at the peak was much dimmer than the upper limit from H.E.S.S. observations. However, we show that either a smaller viewing angle or a larger external density would make similar off-axis events detectable in the future at VHE, even above 100 Mpc with the sensitivity of the CTA. Large external densities are expected in the case of fast mergers, but the existence of a formation channel for such binary neutron stars is still uncertain. We highlight that VHE afterglow detections would help probing efficiently such systems.
@@ 1 Introduction
The first detection of gravitational waves (GW) from a Binary Neutron Star (BNS) merger, GW 170817 (Abbott et al., 2017a), was followed by several electromagnetic counterparts (Abbott et al. 2017b and references therein). A short Gamma-Ray Burst (GRB), GRB 170817A, was detected $\sim 1.7\leavevmode\nobreak\ \mathrm{s}$ after the GW signal (Goldstein et al., 2017; Savchenko et al., 2017); a fast-decaying thermal transient in the visible/infrared range, the kilonova, was observed for $\sim 10$ days after the merger (see e.g. Villar et al., 2017; Tanvir et al., 2017; Cowperthwaite et al., 2017); and a non-thermal afterglow was observed from radio to X-rays for more than three years (see e.g. Balasubramanian et al., 2021; Troja et al., 2020; Hajela et al., 2019). Among the many advances made possible by this exceptional multi-messenger event, GW 170817/GRB 170817A is in particular the first direct association between a BNS merger and a short GRB and has helped to better understand the relativistic ejection associated with such events. Indeed, the study of the relativistic ejecta benefited not only from an exceptional multi-wavelength follow-up, but also from observing conditions very different from those for other short GRBs: a much smaller distance (GW 170817 was hosted in NGC 4993 at $\sim 40\leavevmode\nobreak\ \mathrm{Mpc}$, Palmese et al. 2017; Cantiello et al. 2018) and a significantly off-axis observation ($32^{+10}_{-13}\pm 1.7$ deg derived from the GW signal using the accurate localization and distance of NGC 4993, Finstad et al. 2018).
@@
The prompt short GRB is puzzling: it is extremely weak despite the short distance, but its peak energy is above $150\,\mathrm{keV}$ (Goldstein et al., 2017). It is very unlikely that it is produced by internal dissipation in the ultra-relativistic core jet like in other GRBs at cosmological distance (Matsumoto et al., 2019). GRB 170817A was rather emitted in mildly-relativistic/mildly-energetic material on the line-of-sight: a promising mechanism is the shock-breakout emission when the relativistic core jet emerges from the kilonova ejecta (Bromberg et al., 2018; Gottlieb et al., 2018).
@@
The interpretation of the afterglow is better understood. The slow rise of the light curve until its peak after $\sim 120-160$ days hints towards the off-axis observation of a decelerating jet surrounded by a lateral structure (see e.g. D’Avanzo et al., 2018), possibly inherited from the early interaction with the kilonova ejecta also invoked for the weak prompt emission (Gottlieb et al., 2018). The presence of a relativistic ejecta seen off-axis is confirmed by the compactness of the source and its apparent superluminal motion measured by Very Long Baseline Interferometry (VLBI) imagery (Mooley et al., 2018; Ghirlanda et al., 2019; Mooley et al., 2022). The light curve rise is dominated by the lateral structure of the relativistic ejecta, while its peak and decay are dominated by the deceleration of the ultra-relativistic core jet, which requires a kinetic energy comparable to usual values found in short GRBs (see e.g. Ghirlanda et al., 2019).
@@
Such a lateral structure may be a common feature in GRBs due to the early propagation of the relativistic ejecta through the infalling envelope of the stellar progenitor (long GRBs) or the post-merger ejecta responsible for the kilonova emission (short GRBs) (Bromberg et al., 2011). Possible signatures of such a lateral structure in GRBs at cosmological distance viewed slightly off-axis have been recently discussed, for instance to explain the complex phenomenology observed in the early afterglow (Beniamini et al., 2020a; Oganesyan et al., 2020; Ascenzi et al., 2020; Duque et al., 2022) or the non-standard decay of the afterglow of the extremely bright GRB 221009A (O’Connor et al., 2023; Gill & Granot, 2023). Accounting for the lateral structure of the jet in afterglow models has thus become necessary not only for GW 170817 but for other cosmic GRBs as well.
@@
Most models of the off-axis afterglow of GW 170817 only consider synchrotron emission and are therefore limited to the spectral range of observations, from radio to X-rays. However, upper limits on the flux at very high energy (VHE) were obtained by H.E.S.S. at two epochs, a few days after the merger (Abdalla et al., 2017) and around its peak (Abdalla et al., 2020). Less constraining upper limits were also obtained by HAWC at early times (Galván et al., 2019) and MAGIC around the afterglow peak (Salafia et al., 2022a). Discussing the constraints associated to these upper limits requires to include the Synchrotron Self-Compton (SSC) emission in afterglow models, i.e.  the Inverse Compton (IC) scatterings of the synchrotron photons by the relativistic electrons emitting them. Indeed, recent detections of the VHE afterglow of several long GRBs [^1] Candidate VHE photons are associated at a lower level of confidence with several other bursts by MAGIC, including the short GRB 160821B (Acciari et al., 2021). by H.E.S.S. (GRB 180720B, Abdalla et al. 2019; GRB 190829A, H. E. S. S. Collaboration et al. 2021), MAGIC (GRB 190114C, MAGIC Collaboration et al. 2019a, b; GRB 201216C, Blanch et al. 2020) and LHAASO (GRB 221009A, LHAASO Collaboration 2023) are challenging pure synchrotron afterglow models (see however H. E. S. S. Collaboration et al., 2021) and suggests the dominant contribution of a new emission process at VHE (see the discussion of possible processes at VHE by Gill & Granot, 2022), most probably SSC emission, as suggested for instance by the modelling of the afterglows of GRB 190114C (MAGIC Collaboration et al., 2019b; Wang et al., 2019; Derishev & Piran, 2021), GRB 190829A (Salafia et al., 2022b) or GRB 221009A (Sato et al., 2023).
@@
While SSC emission can be efficiently computed analytically in the Thomson regime (Panaitescu & Kumar, 2000; Sari & Esin, 2001), several studies have pointed out the importance to also account for the Klein-Nishina (KN) attenuation at high-energy (see e.g. Nakar et al., 2009; Murase et al., 2011; Beniamini et al., 2015; Jacovich et al., 2021; Yamasaki & Piran, 2022). In the KN regime, the IC power of an electron strongly depends on its Lorentz factor. This affects the cooling of the electron distribution and therefore also impacts the synchrotron spectrum (Derishev et al., 2001; Nakar et al., 2009; Bošnjak et al., 2009), which can significantly differ from the standard prediction given by Sari et al. (1998) for the pure synchrotron case.
@@
Motivated by these recent advances in GRB afterglow studies, we propose in this paper a consistent model of GRB afterglow emission from a decelerating laterally-structured jet where electrons accelerated at the external forward shock radiate at all wavelengths by synchrotron emission and SSC diffusion. The treatment of SSC in the KN regime follows the approach proposed by Nakar et al. (2009), extended to account for the maximum Lorentz factor of accelerated electrons, and the attenuation at high-energy due to pair creation. The numerical implementation of the model is optimized to be computationally-efficient. This allows to use Bayesian statistics for parameter inference. We then apply this model to the multi-wavelength observations of the afterglow of GW 170817 to understand the physical constraints brought by the H.E.S.S. upper limits and to discuss if the detection of post-merger VHE afterglows may become possible in the future, especially in the coming era of the Cerenkov Telescope Array (CTA, see Cherenkov Telescope Array Consortium et al. 2019).
@@
The model and its assumptions are detailed in Section 2 and then tested and compared to other afterglow models in Section 3. In Section 4 we show the results obtained when fitting GW 170817’s afterglow. We discuss the predicted VHE emission in light of the H.E.S.S. upper limit. We study the conditions for future post-merger detections of VHE afterglows in Section 5 and highlight that VHE emission is favored by a larger external density compared to GW 170817, which may help probing the population of fast-merging binaries, if it exists. Our conclusions are summarized in Section 6.
@@ 2 Modelling the VHE afterglow of structured jets
Our aim is to model the GRB afterglow at all wavelengths from radio bands to the TeV range. We limit ourselves to the contribution of the forward external shock propagating in the external medium and leave to a future work the extension of the model to include the reverse shock contribution at early times. Motivated by the observations of GW 170817, we rather focus here on two other aspects: the lateral structure of the jet and the detailed calculation of the Inverse Compton spectral component, taking into account Klein-Nishina effects.
In this section we describe our model: assumptions for the structure of the initial relativistic outflow (§ 2.1), dynamics of its deceleration (§ 2.2), assumptions for the acceleration of electrons and the amplification of the magnetic field at the shock front (§ 2.3), detailed calculation of the emission in the comoving frame of the shocked external medium (including synchrotron and Inverse Compton components, § 2.4) and finally integration over equal-arrival time surfaces to compute the observed flux in the observer frame (light curves and spectra, § 2.5). Our numerical implementation of this model optimizes the computation time (typically a few seconds to compute light curves and/or spectra at different frequencies on a laptop) so that we can explore the parameter space with a Bayesian approach for data fitting, as described in Section 4. In the following, in case of ambiguity, a physical quantity $q$ is written $q$ without a prime in the fixed source frame, and $q^{\prime}$ with a prime in the comoving frame of the emitting material.
@@ 2.1 Relativistic outflow: geometry and structure
We consider a laterally-structured jet. The initial energy per solid angle $\epsilon_{0}(\theta)$ and the initial Lorentz factor $\Gamma_{0}(\theta)$ decrease with $\theta$, the angle from the jet axis. We define $\theta_{\mathrm{c}}$ as the opening angle of the core, i.e. the ultra-relativistic/ultra-energetic central part of the jet. We then express
@@
where $\epsilon_{0}^{\mathrm{c}}$ is the initial energy per solid angle and $\Gamma_{0}^{\rm c}$ is the initial Lorentz factor, both on the jet axis ($\theta=0$). $f_{a}$ and $g_{b}$ are the normalized profiles for the energy and Lorentz factor in the lateral structure.
Different lateral structures have been suggested in the literature, following the photometric and VLBI observations of GW 170817’s GRB afterglow. In this paper, we model this afterglow assuming the power-law structure used by Duque et al. (2019):
@@
This prescription allows a direct comparison with a top-hat jet ($f_{a}(x)=g_{b}(x)=1$ for $x\leq 1$ and 00 otherwise). For comparison with other studies (see Section 3), we also considered the following possible structures:
  • Power-law jet from Gill & Granot (2018): $f_{a}(x)=X^{-a}$ and $g_{b}(x)=X^{-b}$, with $X=\sqrt{1+x^{2}}$.
  • Power-law jet as defined in afterglowpy, from Ryan et al. (2020): $f_{a}\left(x\right)=\left(\sqrt{1+{x^{2}}/{a}}\right)^{-a}$. The model used in afterglowpy is limited to the self-similar evolution of the jet, which is independent of the initial value of the Lorentz factor. Therefore $g_{b}$ is not specified in this case.
  • Gaussian jet from Gill & Granot (2018): $f_{a}(x)=g_{b}(x)=\max{\left(e^{-{x^{2}}/{2}}\,;e^{-{x_{\mathrm{max}}^{2}}/{2}}\right)}\,,$ where $x_{\mathrm{max}}=\theta_{\mathrm{max}}/\theta_{\mathrm{c}}$ and $\theta_{\rm max}$ is defined as the angle where $\beta_{0,\mathrm{min}}\left(\theta_{\rm max}\right)=0.01$ (by default) or any other specific value. We note that Ryan et al. (2020) use a similar parametrization of the lateral structure for Gaussian jets in afterglowpy, with $\theta_{\rm max}=90\leavevmode\nobreak\ \mathrm{deg}$.
@@
The factor $2$ accounts for the counter-jet. The usual isotropic-equivalent energy of the core jet equals
@@
For the power-law structure considered here (Equation (5)), we have $E_{0}^{\mathrm{c}}=4\pi\epsilon_{0}^{\mathrm{c}}\left(1-\cos{\theta_{\mathrm{c}}}\right)$ and $E_{0,\mathrm{iso}}^{\mathrm{c}}=4\pi\epsilon_{0}^{\mathrm{c}}$.
We do not consider in this paper a possible radial structure of the outflow that may also very well be present due to the variability of the central engine. This radial structure may be at least partially smoothed out during the early propagation, e.g. via internal shocks (Kobayashi et al., 1997; Daigne & Mochkovitch, 1998, 2000) and should have more impact on the reverse shock (see e.g. Uhm & Beloborodov, 2007; Genet et al., 2007), not included in the present version of the model. Therefore $E_{0\rm,iso}^{\rm c}$, $\Gamma_{0}^{\rm c}$, $\theta_{\rm c}$, $a$ and $b$ (or $\beta_{0,\rm min}$ for the Gaussian jet) are the only four free parameters needed to fully describe the initial structure of the jet before the deceleration starts.
In our numerical implementation of the model, we suppress the lateral structure above the maximum angle $\theta_{\mathrm{max}}$. This angle is taken as the maximum between $\theta_{\epsilon}$, defined as the angle up to which a fraction $(1-\epsilon)$ of the jet energy is contained, and the viewing angle $\theta_{\mathrm{v}}$ (see § 2.5). We therefore solve for $\theta_{\epsilon}$ the equation
@@
We use $\epsilon=0.01$ in the following. Ryan et al. (2020) introduce a similar parameter $\theta_{\mathrm{W}}$, to minimize computation time. Taking $\theta_{\mathrm{max}}=\max{\left(\theta_{\epsilon};\theta_{\mathrm{v}}\right)}$ allows to keep a precise calculation at early times even for very large viewing angles.
In practice, the lateral structure is discretized in $N+1$ components, $i=0$ being the core jet and $i=1\to N$ being rings at increasing angles in the lateral structure. Each component is defined by $\theta_{\mathrm{min},i}\leq\theta\leq\theta_{\mathrm{max},i}$ with $\theta_{\mathrm{min},0}=0$ and $\theta_{\mathrm{max},0}=\theta_{\mathrm{c}}$ for the core jet; $\theta_{\mathrm{min},i}=\theta_{\mathrm{max},i-1}$ for $i=1\to N$ and $\theta_{\mathrm{max},N}=\theta_{\mathrm{max}}$. Each component is treated independently with fixed limits in latitude (no lateral spreading). The dynamics is computed such that $R_{i}(t)=R(\theta_{i};t)$ and $\Gamma_{i}(t)=\Gamma(\theta_{i};t)$, where $R(\theta,t)$ and $\Gamma(\theta,t)$ are given by the solution for the dynamics of the deceleration discussed below, and where $\theta_{i}=(\theta_{\mathrm{min},i}+\theta_{\mathrm{max},i})/2$ for $i\geq 1$ and $\theta_{0}=0$. We define the successive $\theta_{\mathrm{min},i}$ and $\theta_{\mathrm{max},i}$ such that each component carries an equal amount of energy (except for the core, $i=0$), though a linear increase can also be chosen. Finally, we typically choose a discretization comprised of $N=15$ components, that we find to be a good compromise between model accuracy and computational efficiency.

### v2

# Very high energy afterglow of structured jets: GW 170817 and prospects for future detections

@@
Very high energy afterglow of structured jets:
@@
Clément Pellouin, Very high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detections    Frédéric Daigne, Very high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detectionsVery high energy afterglow of structured jets: GW 170817 and prospects for future detections
@@
We present a complete numerical model of the afterglow of a laterally structured relativistic ejecta from the radio to very high energies (VHE). This includes a self-consistent calculation of the synchrotron radiation, with its maximum frequency, and of synchrotron self-Compton (SSC) scattering that takes the Klein-Nishina regime into account. Attenuation due to pair production is also included. This model is computationally efficient and allows multi-wavelength data fitting. As a validation test, the radiative model was used to fit the broad-band spectrum of GRB 190114C at 90 s up to the TeV range. The full model was then used to fit the afterglow of GW 170817 and predict its VHE emission. We find that the SSC flux at the peak was much dimmer than the upper limit from H.E.S.S. observations. However, we show that either a smaller viewing angle or a higher external density would make similar off-axis events detectable in the future at VHE, even above 100 Mpc with the sensitivity of the Cherenkov telescope array. High external densities are expected in the case of fast mergers, but the existence of a formation channel for these binary neutron stars is still uncertain. We highlight that VHE afterglow detections would help to efficiently probe systems like this.
@@ 1 Introduction
The first detection of gravitational waves (GWs) from a binary neutron star (BNS) merger, GW 170817 (Abbott et al. 2017a), was followed by the detection of several electromagnetic counterparts (Abbott et al. 2017b and references therein). The short gamma-ray burst GRB 170817A, was detected $\sim 1.7\leavevmode\nobreak\ \mathrm{s}$ after the GW signal (Goldstein et al. 2017; Savchenko et al. 2017); a fast-decaying thermal transient in the visible and infrared range, a kilonova, was observed for $\sim 10$ days after the merger (see e.g. Villar et al. 2017; Tanvir et al. 2017; Cowperthwaite et al. 2017); and a non-thermal afterglow was observed from radio to X-rays for more than three years (see e.g. Balasubramanian et al. 2021; Troja et al. 2020; Hajela et al. 2019). Among the many advances made possible by this exceptional multi-messenger event, GW 170817/GRB 170817A is in particular the first direct association between a BNS merger and a short gamma-ray burst (GRB) and it has helped us to better understand the relativistic ejection associated with events like this. Indeed, the study of the relativistic ejecta benefited not only from an exceptional multi-wavelength follow-up, but also from observing conditions that were very different from those for other short GRBs: a much smaller distance (GW 170817 was hosted in NGC 4993 at $\sim 40\leavevmode\nobreak\ \mathrm{Mpc}$; Palmese et al. 2017; Cantiello et al. 2018), and a significantly off-axis observation ($32^{+10}_{-13}\pm 1.7$ deg derived from the GW signal using the accurate localisation and distance of NGC 4993, Finstad et al. 2018).
@@
The prompt short GRB is puzzling: it is extremely weak despite the short distance, but its peak energy is above $150\,\mathrm{keV}$ (Goldstein et al. 2017). It is very unlikely that it is produced by internal dissipation in the ultra-relativistic core jet like in other GRBs at cosmological distance (Matsumoto et al. 2019). GRB 170817A was rather emitted in mildly relativistic and mildly energetic material on the line of sight: The shock-breakout emission when the relativistic core jet emerges from the kilonova ejecta is a promising mechanism (Bromberg et al. 2018; Gottlieb et al. 2018).
@@
The interpretation of the afterglow is better understood. The slow rise of the light curve until its peak after $\sim 120-160$ days indicates that this was an off-axis observation of a decelerating jet surrounded by a lateral structure (see e.g. D’Avanzo et al. 2018) that may have been inherited from the early interaction with the kilonova ejecta. This was also invoked for the weak prompt emission (Gottlieb et al. 2018). The relativistic ejecta seen off-axis is also confirmed by the compactness of the source and its apparent superluminal motion as measured by very long baseline interferometry (VLBI) imagery (Mooley et al. 2018; Ghirlanda et al. 2019; Mooley et al. 2022). The light curve rise is dominated by the lateral structure of the relativistic ejecta, while its peak and decay are dominated by the deceleration of the ultra-relativistic jet core, which requires a kinetic energy comparable to values commonly found in short GRBs (see e.g. Ghirlanda et al. 2019).
@@
A lateral structure like this may be a common feature in GRBs due to the early propagation of the relativistic ejecta, either through the infalling envelope of the stellar progenitor (long GRBs), or through the post-merger ejecta that produces the kilonova emission (short GRBs) (Bromberg et al. 2011). Possible signatures of this lateral structure in GRBs at cosmological distance viewed slightly off-axis were recently discussed, for instance to explain the complex phenomenology observed in the early afterglow (Beniamini et al. 2020a; Oganesyan et al. 2020; Ascenzi et al. 2020; Duque et al. 2022) or the non-standard decay of the afterglow of the extremely bright GRB 221009A (O’Connor et al. 2023; Gill & Granot 2023). Accounting for the lateral structure of the jet in afterglow models has thus become necessary not only for GW 170817, but for other cosmic GRBs as well.
@@
Most models of the off-axis afterglow of GW 170817 only considered synchrotron emission and are therefore limited to the spectral range of observations, from radio to X-rays. However, upper limits on the flux at very high energy (VHE) were obtained by the high energy stereoscopic system telescopes (H.E.S.S.) at two epochs, a few days after the merger (Abdalla et al. 2017) and around its peak (Abdalla et al. 2020). Less constraining upper limits were also obtained by the high-altitude water Cherenkov gamma-ray observatory (HAWC) at early times (Galván et al. 2019) and by the major atmospheric gamma-ray imaging Cherenkov telescope (MAGIC) around the afterglow peak (Salafia et al. 2022a). A discussion of the constraints associated with these upper limits requires that the synchrotron self-Compton (SSC) emission is included in afterglow models, that is, the inverse Compton (IC) scattering of the synchrotron photons by the relativistic electrons that emit them. Recent detections of the VHE afterglow of several long GRBs [^1] Candidate VHE photons are associated at a lower level of confidence with several other bursts by MAGIC, including the short GRB 160821B (Acciari et al. 2021). by H.E.S.S. (GRB 180720B; Abdalla et al. 2019; GRB 190829A; H. E. S. S. Collaboration et al. 2021), MAGIC (GRB 190114C; MAGIC Collaboration et al. 2019a, b; GRB 201216C; Blanch et al. 2020), and the large high altitude air shower observatory (LHAASO) (GRB 221009A; LHAASO Collaboration 2023) challenge purely synchrotron afterglow models (see however H. E. S. S. Collaboration et al. 2021) and suggest the dominant contribution of a new emission process at VHE (see the discussion of possible processes at VHE by Gill & Granot 2022), most probably, SSC emission, as suggested for instance by the modelling of the afterglows of GRB 190114C (MAGIC Collaboration et al. 2019b; Wang et al. 2019; Derishev & Piran 2021), GRB 190829A (Salafia et al. 2022b), or GRB 221009A (Sato et al. 2023).
@@
While SSC emission can be efficiently computed analytically in the Thomson regime (Panaitescu & Kumar 2000; Sari & Esin 2001), several studies have pointed out that the Klein-Nishina (KN) attenuation at high energy also needs to be accounted for (see e.g. Nakar et al. 2009; Murase et al. 2011; Beniamini et al. 2015; Jacovich et al. 2021; Yamasaki & Piran 2022). In the KN regime, the IC power of an electron strongly depends on its Lorentz factor. This affects the cooling of the electron distribution, and it therefore also impacts the synchrotron spectrum (Derishev et al. 2001; Nakar et al. 2009; Bošnjak et al. 2009), which can significantly differ from the standard prediction given by Sari et al. (1998) for the purely synchrotron case.
@@
Motivated by these recent advances in GRB afterglow studies, we propose in this paper a consistent model of GRB afterglow emission from a decelerating laterally structured jet where electrons accelerated at the external forward shock radiate at all wavelengths by synchrotron emission and SSC scattering. The treatment of SSC in the KN regime follows the approach proposed by Nakar et al. (2009), extended to account for the maximum Lorentz factor of the accelerated electrons and the attenuation at high energy due to pair creation. The numerical implementation of the model was optimised to be computationally efficient. This allowed to use Bayesian statistics to infer the parameters. We then applied this model to the multi-wavelength observations of the afterglow of GW 170817 to understand the physical constraints brought by the H.E.S.S. upper limits and to discuss whether the detection of post-merger VHE afterglows may become possible in the future, especially in the coming era of the Cherenkov telescope array (CTA; see Cherenkov Telescope Array Consortium et al. 2019).
@@
The model and its assumptions are detailed in Sect. 2 and are then tested and compared to other afterglow models in Sect. 3. In Sect. 4 we show the results obtained when fitting the afterglow of GW 170817. We discuss the predicted VHE emission in the light of the H.E.S.S. upper limit. We study the conditions for future post-merger detections of VHE afterglows in Sect. 5 and highlight that VHE emission is favoured by a higher external density compared to GW 170817, which may help to probe the population of fast-merging binaries, if it exists. Our conclusions are summarised in Sect. 6.
@@ 2 Modelling the VHE afterglow of structured jets
Our aim is to model the GRB afterglow at all wavelengths from radio bands to the TeV range. We limit ourselves to the contribution of the forward external shock propagating in the external medium and leave the extension of the model to a future work that includes the reverse-shock contribution at early times. Motivated by the observations of GW 170817, we rather focus here on two other aspects: the lateral structure of the jet, and the detailed calculation of the IC spectral component, including KN effects.
In this section, we describe our model: The assumptions for the structure of the initial relativistic outflow (Sect. 2.1), the dynamics of its deceleration (Sect. 2.2), the assumptions for the acceleration of electrons and the amplification of the magnetic field at the shock front (Sect. 2.3), the detailed calculation of the emission in the comoving frame of the shocked external medium (including synchrotron and IC components; Sect. 2.4), and finally, the integration over equal-arrival time surfaces to compute the observed flux in the observer frame (light curves and spectra; Sect. 2.5). Our numerical implementation of this model optimises the computation time (typically a few seconds for computing light curves and/or spectra at different frequencies on a laptop) so that we can explore the parameter space with a Bayesian approach for the data fitting, as described in Sect. 4. In case of ambiguity, a physical quantity $q$ is written $q$ without a prime in the fixed source frame, and $q^{\prime}$ with a prime in the comoving frame of the emitting material.
@@ 2.1 Relativistic outflow: Geometry and structure
We consider a laterally structured jet. The initial energy per solid angle $\epsilon_{0}(\theta)$ and the initial Lorentz factor $\Gamma_{0}(\theta)$ decrease with $\theta$, which is the angle from the jet axis. We define $\theta_{\mathrm{c}}$ as the opening angle of the core, that is, the ultra-relativistic/ultra-energetic central part of the jet. We then write
@@
where $\epsilon_{0}^{\mathrm{c}}$ is the initial energy per solid angle, and $\Gamma_{0}^{\rm c}$ is the initial Lorentz factor, both on the jet axis ($\theta=0$). $f_{a}$ and $g_{b}$ are the normalised profiles for the energy and Lorentz factor in the lateral structure.
Different lateral structures have been suggested in the literature, following the photometric and VLBI observations of the GRB afterglow of GW 170817. In this paper, we model this afterglow assuming the power-law structure used by Duque et al. (2019),
@@
This prescription allows a direct comparison with a top-hat jet ($f_{a}(x)=g_{b}(x)=1$ for $x\leq 1$ and 00 otherwise). For comparison with other studies (see Sect. 3), we also considered the following possible structures:
  • A power-law jet from Gill & Granot (2018), $f_{a}(x)=X^{-a}$ and $g_{b}(x)=X^{-b}$, with $X=\sqrt{1+x^{2}}$.
  • A power-law jet as defined in afterglowpy, from Ryan et al. (2020), $f_{a}\left(x\right)=\left(\sqrt{1+{x^{2}}/{a}}\right)^{-a}$. The model used in afterglowpy is limited to the self-similar evolution of the jet, which is independent of the initial value of the Lorentz factor. Therefore, $g_{b}$ is not specified in this case.
  • A Gaussian jet from Gill & Granot (2018), $f_{a}(x)=g_{b}(x)=\max{\left(e^{-{x^{2}}/{2}}\,;e^{-{x_{\mathrm{max}}^{2}}/{2}}\right)}\,,$ where $x_{\mathrm{max}}=\theta_{\mathrm{max}}/\theta_{\mathrm{c}}$ and $\theta_{\rm max}$ is defined as the angle where $\beta_{0,\mathrm{min}}\left(\theta_{\rm max}\right)=0.01$ (by default) or any other specific value. We note that Ryan et al. (2020) used a similar parametrisation of the lateral structure for Gaussian jets in afterglowpy, with $\theta_{\rm max}=90\leavevmode\nobreak\ \mathrm{deg}$.
@@
The factor of $2$ accounts for the counter-jet. The usual isotropic-equivalent energy of the core jet equals
@@
For the power-law structure considered here (Eq. (5)), we have $E_{0}^{\mathrm{c}}=4\pi\epsilon_{0}^{\mathrm{c}}\left(1-\cos{\theta_{\mathrm{c}}}\right)$ and $E_{0,\mathrm{iso}}^{\mathrm{c}}=4\pi\epsilon_{0}^{\mathrm{c}}$.
We did not consider a possible radial structure of the outflow that may also very well be present due to the variability of the central engine. This radial structure may be at least partially smoothed out during the early propagation, for example via internal shocks (Kobayashi et al. 1997; Daigne & Mochkovitch 1998, 2000), and is expected to have more impact on the reverse shock (see e.g. Uhm & Beloborodov 2007; Genet et al. 2007), which is not included in the current version of the model. Therefore, $E_{0\rm,iso}^{\rm c}$, $\Gamma_{0}^{\rm c}$, $\theta_{\rm c}$, $a$, and $b$ (or $\beta_{0,\rm min}$ for the Gaussian jet) are the only four free parameters needed to fully describe the initial structure of the jet before the deceleration starts.
In our numerical implementation of the model, we suppressed the lateral structure above the maximum angle $\theta_{\mathrm{max}}$. This angle is taken as the maximum between $\theta_{\epsilon}$, defined as the angle up to which a fraction $(1-\epsilon)$ of the jet energy is contained, and the viewing angle $\theta_{\mathrm{v}}$ (see Sect. 2.5). We therefore solved for $\theta_{\epsilon}$ the equation
@@
We use $\epsilon=0.01$ in the following. Ryan et al. (2020) introduced a similar parameter $\theta_{\mathrm{W}}$ to minimise computation time. Taking $\theta_{\mathrm{max}}=\max{\left(\theta_{\epsilon};\theta_{\mathrm{v}}\right)}$ allowed us to keep a precise calculation at early times even for very large viewing angles.
In practice, the lateral structure is discretised into $N+1$ components, $i=0$ being the core jet and $i=1\to N$ being rings at increasing angles in the lateral structure. Each component is defined by $\theta_{\mathrm{min},i}\leq\theta\leq\theta_{\mathrm{max},i}$ with $\theta_{\mathrm{min},0}=0$ and $\theta_{\mathrm{max},0}=\theta_{\mathrm{c}}$ for the core jet; $\theta_{\mathrm{min},i}=\theta_{\mathrm{max},i-1}$ for $i=1\to N$, and $\theta_{\mathrm{max},N}=\theta_{\mathrm{max}}$. Each component is treated independently with fixed limits in latitude (no lateral spreading). The dynamics is computed such that $R_{i}(t)=R(\theta_{i};t)$ and $\Gamma_{i}(t)=\Gamma(\theta_{i};t)$, where $R(\theta,t)$ and $\Gamma(\theta,t)$ are given by the solution for the dynamics of the deceleration discussed below, and where $\theta_{i}=(\theta_{\mathrm{min},i}+\theta_{\mathrm{max},i})/2$ for $i\geq 1$ and $\theta_{0}=0$. We define the successive $\theta_{\mathrm{min},i}$ and $\theta_{\mathrm{max},i}$ such that each component carries an equal amount of energy (except for the core, $i=0$), although a linear increase can also be chosen. Finally, we typically choose a discretisation comprised of $N=15$ components, which we find to be a good compromise between model accuracy and computational efficiency.